(1)Portland plc is a soft drink producer which is considering to acquire a food and drink retailer, Starland plc, via a hostile takeover. Financial advisors claim that a combined equity will allow Portland plc to save up to £4 million in the future by bringing the production and distribution of soft drinks under the control of one firm. Both companies are solely financed by equity. Portland plc has 17 million shares at a price of £5.50 each, while Starland plc has 6 million shares at a price of £8.20 each. Portland’s top executives suggest that an initial bid with a premium of 32 percent is ample to persuade the shareholder of the target company to sell their holdings. Portland plc has enough cash reserves to fund the takeover bid. You are asked to assess the proposed takeover from the point of view of Portland’s shareholders assuming that the cost of capital of the combined entity is 20 percent.

(2)Suppose that Portland plc intends to bid for Starland plc using equity instead of cash. The company has been advised to offer 2 shares of Portland plc for every 1 share of Starland plc. What is the percentage premium offered to the shareholders of Starland plc? Assess the merger from the point of view of Portland’s shareholders.
Answer #1
1.NPV of the proposed takeover from Portland shareholders' point of view= PV of savings -Purchase cost of Starland's shares at a premium of 32%, ie.
(4000000/20%)-((6000000*8.2)*1.32)=
-44944000
2. Percentage premium offered= at the rate of 2 shares of Portland to 1 share of Starland----- ie.
((5.5*2)-8.2)/8.2=
34.15%
NPV of the merger from Portland shareholders' point of view=Just the PV of savings   & allotment of shares at a premium of 34.15%
ie. 4000000/20%=
20000000
(now for all the 17 mln.+(6*2=)12 mln.=29 mln. Shareholders.